OpenStudy (anonymous):

(x-4)^2/36 + (y+2)^2/16 =1 What is the sum of the focal radii? What are the coordinates of the foci?

OpenStudy (anonymous):

\[\frac{ (x-4)^2 }{ 36 } + \frac{ (y+2)^2 }{ 16 } =1\]

OpenStudy (anonymous):

help please

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-4)^2}{36}+\cfrac{(y+2)^2}{16}=1\implies \cfrac{(x-4)^2}{6^2}+\cfrac{(y+2)^2}{4^2}=1\\ \quad \\ \implies \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\\ \quad \\ \quad \\ \textit{distance from the vertex to the foci is "c"}\quad c=\sqrt{a^2-b^2}\)

OpenStudy (anonymous):

is that the answer to the first question

OpenStudy (jdoe0001):

hmmm I assume by "focal radii" it's meant the sum of both distances from the vertex

OpenStudy (anonymous):

I Have no clue honestly.. My teacher just gave me work and didn't even give me a lesson over this.

OpenStudy (jdoe0001):

so all you really need to find is the "c" distance, to get the sum of them and their coordinates, notice that the "a" component is the BIGGER denominator and is under the "x" variable, thus that means the ellipse is of traverse horizontal axis, or a horizontal one

OpenStudy (anonymous):

how do you find the coordinates? -_-

OpenStudy (jdoe0001):

well, what did you get for "c"?

OpenStudy (anonymous):

2?

OpenStudy (jdoe0001):

2?

OpenStudy (anonymous):

6-4 = 2..

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-4)^2}{\color{red}{36}}+\cfrac{(y+2)^2}{\color{red}{16}}=1\implies \cfrac{(x-4)^2}{6^2}+\cfrac{(y+2)^2}{4^2}=1\\ \quad \\ \implies \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\\ \quad \\ \quad \\ \textit{distance from the vertex to the foci is "c"}\quad c=\sqrt{\color{red}{a^2-b^2}}\)

OpenStudy (jdoe0001):

notice, is the squared version of "a" and "b"

OpenStudy (anonymous):

so 10? .. i dont know what im doing.. i thought 2 was right

OpenStudy (jdoe0001):

well, a = 6, b = 4 \(\bf a^2= 36\quad b^2= 16\)

OpenStudy (jdoe0001):

\(\bf \cfrac{(x-4)^2}{36}+\cfrac{(y+2)^2}{16}=1\\ \quad \\ c=\sqrt{a^2-b^2}\implies c=\sqrt{20}\implies c=2\sqrt{5}\)

OpenStudy (anonymous):

wow.. im dumb.. lol thanks

OpenStudy (jdoe0001):

notice the center of the parabola, at (h, k) \(\bf \cfrac{(x-4)^2}{36}+\cfrac{(y+2)^2}{16}=1\implies \cfrac{(x-(\color{red}{+4}))^2}{36}+\cfrac{(y-(\color{red}{-2}))^2}{16}=1\\ \quad \\ \implies \cfrac{(x-h)^2}{a^2}+\cfrac{(y-k)^2}{b^2}=1\) so your center is at ( 4, -2 ) is a horizontal ellipse, like  so the focii are at \(\bf (4 \pm 2\sqrt{5}\quad ,\quad -2)\)

OpenStudy (jdoe0001):

and their sum would be \(\bf 2\sqrt{5}+2\sqrt{5}\) I'd think

OpenStudy (anonymous):

thanks :) i also have to find the eccentricity of the ellipse.

OpenStudy (jdoe0001):

\(\bf \textit{eccentricity of an ellipse}=\cfrac{c}{a}\)

OpenStudy (anonymous):

oh okay gracias

OpenStudy (jdoe0001):

yw